# Asymptotic Consistency and Generalization in Hybrid Models of Regularized Selection and Nonlinear Learning

### v1

@@ Abstract
This study explores how different types of supervised models perform in the task of predicting and selecting relevant variables in high-dimensional contexts, especially when the data is very noisy. We analyzed three approaches: regularized models (such as Lasso, Ridge, and Elastic Net), black-box models (such as Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and hybrid models that combine both approaches: regularization with nonlinear algorithms. Based on simulations inspired by the Friedman equation, we evaluated 23 models using three complementary metrics: RMSE, Jaccard index, and recall rate. The results reveal that, although black-box models excel in predictive accuracy, they lack interpretability and simplicity, essential factors in many real-world contexts. Regularized models, on the other hand, proved to be more sensitive to an excess of irrelevant variables. In this scenario, hybrid models stood out for their balance: they maintain good predictive performance, identify relevant variables more consistently, and offer greater robustness, especially as the sample size increases. Therefore, we recommend using this hybrid framework in market applications, where it is essential that the results make sense in a practical context and support decisions with confidence.
@@ 1  Introduction
Predictive models for high-dimensional problems pose a significant computational challenge. In many cases, the direct use of nonlinear algorithms to reduce the number of features can be computationally expensive. Although widely used, this framework can generate biases, overfitting (induced by heuristics), and limitations in methods based on variable importance, given that the predictive model itself influences the prioritization of selected variables [1, 2]. The use of non-heuristic methods, such as regularizations independent of the predictive model, has gained prominence in several fields to mitigate such biases, with hybrid models being a promising solution [3]. However, these associated models require further theoretical development regarding their generalization capacity and asymptotic properties.
@@
This work attempts to develop a theoretical framework that analyzes the selection consistency and generalization limits of these hybrid models, focusing on the combination of regularized methods such as Lasso, Ridge, and Elastic Net with widely used black-box algorithms, including Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM.
@@
Regarding the selection step using regularized models, Lasso ([4, 5]) is a selection and regularization method that promotes sparsity in the coefficients through the penalty $\ell 1$. Lasso’s consistency in recovering the true set of variables depends on conditions such as unrepresentability and strong signal strength, which are essential in high-dimensional scenarios. Ridge Regression ([6]) applies quadratic penalty ($\ell 2$) and does not perform variable selection, but contributes to greater stability and variance reduction in the presence of multicollinearity. Elastic Net [7] combines the penalties $\ell 1$ and $\ell 2$, offering a robust alternative to Lasso, especially effective in the face of high correlation between predictors.
@@
Black-box models are widely used in supervised learning applications due to their high predictive capacity and flexibility. Random Forest is based on bagging [8], a technique that reduces variance by combining decision trees generated on bootstrapped subsets. Boosting-based models, such as XGBoost, LightGBM, and CatBoost, use sequential additive learning to reduce bias and improve performance, while presenting challenges in terms of stability and overfitting [9]. H2O GBM offers a scalable implementation of boosting with automatic hyperparameter adjustments. The theory underlying these models includes interpretations such as adaptive neighbors for Random Forest [10] and bias-variance analysis for more randomized ensembles such as extra-trees [11].
@@
Algorithmic stability plays a central role in theoretical analysis and is directly connected to the behavior of generalization errors, as discussed by Bousquet and Elisseeff. In this context, the notion of expected risk consistency stands out, related to the Empirical Risk Minimization Principle (ERM), as presented by Vapnik. This principle establishes that a good model must present asymptotic convergence between the empirical risk observed in the data and the expected risk associated with the true function that generated the data. Such convergence is essential in supervised learning scenarios, as it ensures that, as the number of observations increases, the model’s performance tends to approach optimality within the class of functions considered.
@@
This property is particularly relevant in scenarios where predictive accuracy is more important than the exact recovery of relevant variables, but here it is used to demonstrate the assertive power of the proposed hybrid models. In this context, even without identifying all the true support, a model can be considered predictive consistent if its predictive ability remains robust as the number of observations increases. Formal discussions of this form of consistency can be found in Vapnik, Bousquet and Elisseeff, and Bousquet and Elisseeff. The loss function and the type of penalty also impact the bias and variance of the estimator, Rosasco2004.
@@
To empirically evaluate the proposed framework, controlled simulations were performed based on the Friedman equation, Friedman1991. This equation is widely used in the literature to generate synthetic data with a known ground truth structure, including relevant predictor variables and noise components. The experimental design allows a clear distinction between the true model and the estimated model, enabling the evaluation of method performance in different dimensionality regimes (n and p) and under different levels of noise and structural complexity of the data. Despite the vast literature on regularized methods and black-box models individually, there is a lack of a theoretical framework that rigorously connects and grounds these steps in the hybrid context. This article aims to contribute to filling this gap by offering new theoretical constraints and conditions for the consistency and generalization of the proposed hybrid pipeline.
@@ 2  Methodology
Algorithm 1 Simulation and Evaluation Framework with Regularization and Hybrid Models
1:Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathbb{R}$; sample sizes $n\in\{200,500,1000\}$; number of predictors $p\in\{5,10,50\}$; replicates $n_{\text{sim}}$; performance metrics: RMSE, Jaccard index, Recovery.
@@
Note. We set the upper bound of $m$ to $\min(10,p-1)$ to avoid trivial subsets including all predictors, which would artificially inflate support recovery metrics.
@@
Let $S\subset\{1,\dots,p\}$ be the true set of relevant variables and $\hat{S}$ be the set estimated by the model. We say that a variable selection method is selection consistent if
$P\left(\hat{S}=S\right)\xrightarrow[n\to\infty]{}=1.$ (1)
@@
$\mathbb{E}\left[\left(f(\mathbf{X})-\hat{f}(\mathbf{X})\right)^{2}\right]\xrightarrow[n\to\infty]{}=0,$ (2)
@@
Consistency Proposition of the Hybrid Framework.
Suppose that the true model has a sparse structure with support $S\subset\{1,\dots,p\}$, and that the regularized selection step satisfies the conditions of unrepresentability and signal strength, ensuring $\mathbb{P}(\hat{S}=S)\to 1$ when $n\to\infty$. Further suppose that the predictive model fitted over the selected subspace (black-box model) is consistently predictive in the sense of $R(\hat{f})\to R(f^{*})$. Under these assumptions, the resulting hybrid framework tends to be consistent in both support recovery and predictive risk.
$\mathbb{P}(\hat{S}=S)\to 1\quad\text{e}\quad R(\hat{f})\to R(f^{*})$ (5)
This formulation does not constitute a formal theorem in the literature, but expresses a principle proposed in this work based on two widely recognized theoretical pillars: the consistency of selection via regularization and the predictive consistency of models fitted on relevant subsets [12].
@@ 2.1 Experimental Study
To empirically validate our findings, we chose to conduct simulation studies. The simulations were based on the Friedman model, widely used in the literature to evaluate variable selection methods in nonlinear and high-dimensional scenarios. What is considered a true model corresponds to the case in which the predictor vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$ has real support $S=\{1,2,3,4,5\}$, meaning that only the first five variables affect the response.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ represents the error term, and each covariate $x_{ij}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(0,1),\ for\ j=1,\dots,p$.
Although only the first five variables are present in the response generating function $f(\mathbf{x}_{i})$, the vector $\mathbf{x}_{i}$ contains up to $p=50$ covariates, of which p5p-5 are irrelevant and represent stochastic noise in the model.
@@ 2.2 Models Evaluated
Finally, the third group consists of hybrid models, in which each black-box model was fitted exclusively on the subset of variables previously selected by one of the regularized models, seeking to combine parsimony with predictive flexibility.
@@ 2.3 Simulations
This study evaluated the performance of regularized, pure, and hybrid black-box models under twelve distinct scenarios, defined by combinations of three sample sizes ($n\in\{200,500,1000\}$) and four dimensionality levels ($p\in\{5,10,50,100\}$). For each scenario, ten simulated datasets were generated independently, totaling 120 replications in total.
For each replication, 23 models were applied: 3 regularized models (Ridge, Lasso, and Elastic Net), 5 pure black-box models (Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and 15 hybrid models resulting from the combination of regularizers and black-box algorithms.
Thus, each scenario involved 230 runs (10 datasets × 23 models), totaling 2,760 runs throughout the simulation. Table 1 details the scenarios considered and the composition of the models evaluated by type.
@@
For each combination of sample size $n$ and number of predictors $p$, 10 independent simulations were conducted to ensure robustness in the evaluations. In each of these replications, the data were initially randomly partitioned into training (80%) and test (20%) sets.
The regularized models were then adjusted using the penalty corresponding to the optimal value $\lambda_{\text{min}}$, obtained by cross-validation using the lowest RMSE criterion per feature subset. Based on the set of variables selected by these models, the hybrid models were then adjusted, limiting themselves exclusively to the estimated support.
For comparison purposes, the pure black-box models were also adjusted, considering all available predictors, regardless of relevance. The final performance of the models was then evaluated based on the RMSE (Root Mean Squared Error), Jaccard index, and Recovery metrics.
@@
In Section  3, we focus on the $n=\{200,500,1000\}$ values of the main results for summary and clarity, although the full range evaluated is documented in the appendices and supplementary scripts.
@@ 2.5 Computer Implementation
The results were saved in .csv and .RData files for reproducibility. Graphical visualizations were generated with the ggplot2 and viridis packages.
@@ 3  Results
In this section, we will present the results obtained in the simulation studies. The evaluation compared different modeling strategies, including regularized approaches, black-box models, and hybrid models that combine selection via regularization with nonlinear predictors. The evaluation metrics considered were the root mean square error (RMSE), the Jaccard index (for similarity between selected variables and the true set), and the recovery rate of relevant variables. The models in the subsections are analyzed segmented, allowing us to evaluate the relative performance of each class of approach under different complexity and generalization criteria.
@@ 3.1 Pure Black-Box Models
Table 2 presents the results of the main black-box models in their pure form, denoted as Full because they use the same algorithm for both variable selection and prediction, without combining them with regularized methods. Five algorithms are considered: CatBoost, H2O GBM, LightGBM, Random Forest, and XGBoost. The comparison includes three metrics at different sample sizes: root mean square error (RMSE), Jaccard index, and recall rate.
Table 2: Best performing pure black-box models with 95% noise ($p=50$).
@@
CatBoost_Full RMSE 2.02 (0.22) 1.68 (0.26) 1.51 (0.29)
CatBoost_Full Jaccard 0.85 (0.13) 0.89 (0.11) 0.93 (0.10)
CatBoost_Full Recovery 0.94 (0.09) 0.95 (0.09) 0.95 (0.09)
H2OGBM_Full RMSE 2.13 (0.20) 1.90 (0.19) 1.72 (0.21)
H2OGBM_Full Jaccard 0.88 (0.13) 0.83 (0.14) 0.85 (0.14)
H2OGBM_Full Recovery 0.95 (0.09) 0.95 (0.09) 0.95 (0.09)
LightGBM_Full RMSE 2.16 (0.19) 1.86 (0.21) 1.68 (0.23)
LightGBM_Full Jaccard 0.78 (0.16) 0.80 (0.13) 0.89 (0.12)
LightGBM_Full Recovery 0.95 (0.09) 0.95 (0.09) 0.95 (0.09)
RF_Full RMSE 2.52 (0.13) 2.21 (0.13) 1.99 (0.14)
RF_Full Jaccard 0.80 (0.07) 0.82 (0.01) 0.82 (0.01)
RF_Full Recovery 0.94 (0.10) 0.95 (0.09) 0.95 (0.09)
XGBoost_Full RMSE 2.33 (0.20) 2.03 (0.19) 1.81 (0.23)
XGBoost_Full Jaccard 0.86 (0.12) 0.88 (0.11) 0.92 (0.10)
XGBOOST _FULL RECOVERY  TEXTBF 0.94 (0.09)  TextBF 0.95 (0.09)  TextBF 0.95 (0.09)
The results demonstrated that black-box models stand out for their high predictive power, with RMSE values consistently lower than those of the other approaches. The CatBoost_Full model obtained the lowest errors across all sample sizes, with $RMSE=1.51$ when $n=1000$ standing out. It also presented high Jaccard and Recovery values, indicating stability in identifying relevant variables. In terms of overlap with the true support, the XGBoost_Full, H2OGBM_Full, and LightGBM_Full models also demonstrated excellent performance, with recovery rates equal to or close to $0.95$ in all scenarios. The Jaccard index, in turn, was also high for H2OGBM_Full ($J=0.88$ at $n=200$) and for XGBoost_Full ($J=0.92$ at $n=1000$).

### v2

# Asymptotic Consistency and Generalization in Hybrid Algorithms of Regularized Selection and Nonlinear Learning

@@ Abstract
This study explores how different types of supervised models perform in the task of predicting and selecting relevant variables in high-dimensional contexts, especially when the data is very noisy. We analyzed three approaches: regularized models (Lasso, Ridge, and Elastic Net), black-box models (Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and hybrid Algorithms that combine both approaches: regularization with nonlinear algorithms. Using simulations based on the Friedman equation, we evaluated 23 algorithms using three complementary metrics: RMSE, Jaccard index, and recovery rate. The results revealed that, although black-box models excel in predictive accuracy, they lack interpretability and simplicity, essential factors in many real-world contexts. Regularized models, on the other hand, proved to be more sensitive to an excess of irrelevant variables. In this scenario, hybrid Algorithms stood out for their balance: they maintain good predictive performance, identify relevant variables more consistently, and offer greater robustness, especially as the sample size increases. Therefore, we recommend using this hybrid framework in market applications, where it is essential that the results make sense in a practical context and support decisions with confidence.
@@ 1  Introduction
Predictive models for high-dimensional problems pose a significant computational challenge. In many cases, the direct use of nonlinear algorithms to reduce the number of features can be computationally expensive. Although widely used, this framework can generate biases, overfitting (induced by heuristics), and limitations in methods based on variable importance, given that the predictive model itself influences the prioritization of selected variables [1, 2]. The use of non-heuristic methods, such as regularization independent of the predictive model, has gained prominence in several fields to mitigate such biases, with hybrid Algorithms being a promising solution [3]. However, these associated models require further theoretical development regarding their generalization capacity and asymptotic properties.
@@
This work attempts to develop a theoretical framework that analyzes the selection consistency and generalization limits of these hybrid Algorithms, focusing on the combination of regularized methods such as Lasso, Ridge, and Elastic Net with widely used black-box algorithms, including Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM.
@@
Regarding the selection step using regularized models, Lasso ([4, 5]) is a selection and regularization method that promotes sparsity in the coefficients through the $\ell_{1}$ penalty. Lasso is consistent in recovering the true set of variables depends on conditions such as unrepresentability and strong signal strength, which are essential in high-dimensional scenarios. Ridge Regression ([6]) applies the quadratic penalty ($\ell_{2}$) and does not perform variable selection, but contributes to greater stability and variance reduction in the presence of multicollinearity. Elastic Net [7] combines the $\ell_{1}$ and $\ell_{2}$ penalties, offering a robust alternative to Lasso, which is especially effective in the presence of high correlations between predictors.
@@
Black-box models are widely used in supervised learning applications due to their high predictive capacity and flexibility. Random Forest is based on bagging [8], a technique that reduces variance by combining decision trees generated on bootstrapped subsets. Boosting-based models, such as XGBoost, LightGBM, and CatBoost, use sequential additive learning to reduce bias and improve performance, while presenting challenges in terms of stability and overfitting [9]. H2O GBM offers a scalable implementation of boosting with automatic hyperparameter adjustments. The theory underlying these models includes interpretations such as adaptive neighbors for Random Forest [10] and bias-variance analysis for more randomized ensembles such as Extremely Randomized Trees[11].
@@
Algorithmic stability plays a central role in theoretical analysis and is directly connected to the behavior of generalization errors, as discussed by Bousquet and Elisseeff (2002) [12]. In this context, the notion of expected risk consistency stands out. This is related to the Empirical Risk Minimization Principle (ERM), as presented by Vapnik [13](1999). This principle establishes that a good model must present asymptotic convergence between the empirical risk observed in the data and the expected risk associated with the true function that generated the data. Such convergence is essential in supervised learning scenarios, as it ensures that, as the number of observations increases, the model’s performance tends to approach optimality within the class of functions considered.
@@
This property is particularly relevant in scenarios where predictive accuracy is more important than the exact recovery of relevant variables, but here it is used to demonstrate the assertive power of the proposed hybrid models. In this context, even without identifying all the true support, a model can be considered predictive consistent if its predictive ability remains robust as the number of observations increases. The loss function and the type of penalty also impact the bias and variance of the estimator, [14].
@@
To empirically evaluate the proposed framework, controlled simulations were performed based on the Friedman equation, Friedman (1991) [15]. This equation is widely used in the literature to generate synthetic data with a known ground truth structure, including relevant predictor variables and noise components. The experimental design allows a clear distinction between the true model and the estimated model, enabling the evaluation of method performance in different dimensionality regimes ($n$ and $p$) and under different levels of noise and structural complexity of the data. Despite the vast literature on regularized methods and black-box models used individually, there is a lack of a theoretical framework that rigorously connects and grounds these steps in the hybrid context. This article aims to contribute to filling this gap by offering new theoretical constraints and conditions for the consistency and generalization of the proposed hybrid pipeline.
@@ 2  Methodology
Algorithm 1 Simulation and Evaluation Framework with Hybrid Algorithms
1:Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathbb{R}$; sample sizes $n\in\{50,100,200,500,1000\}$; number of predictors $p\in\{5,10,50,100\}$; replicates $n_{\text{sim}}$; performance metrics: RMSE, Jaccard index, Recovery.
@@
To characterize methods that correctly identify the truly relevant variables in the asymptotic limit, following [16] and [17], one can define $S\subset\{1,\dots,p\}$ as the truly relevant set of variables and $\hat{S}$ as the set estimated by the model. We say that a variable selection method is selection-consistent if
$P\left(\hat{S}=S\right)\xrightarrow[n\to\infty]{}1.$ (1)
@@
$\mathbb{E}\left[\left(f(\mathbf{X})-\hat{f}(\mathbf{X})\right)^{2}\right]\xrightarrow[n\to\infty]{}0,$ (2)
@@
Consistency of the Hybrid Framework.
Suppose that the true model has a sparse structure with support $S\subset\{1,\dots,p\}$, and that the regularized selection step satisfies the conditions of unrepresentability and signal strength, ensuring $\mathbb{P}(\hat{S}=S)\to 1$ when $n\to\infty$. Further suppose that the predictive model fitted over the selected subspace (black-box model) is consistently predictive in the sense of $R(\hat{f})\to R(f^{*})$. Under these assumptions, the resulting hybrid framework would also be consistent in both support recovery and predictive risk.
This conjecture expresses a principle proposed in this work based on two widely recognized theoretical pillars: the consistency of selection via regularization and the predictive consistency of models fitted on relevant subsets [18].
@@ 2.1 Experimental Study
To empirically validate our findings, we chose to carry out simulation studies. The simulations were based on the Friedman model, widely used in the literature to evaluate variable selection methods in nonlinear and high-dimensional scenarios. What is considered a true model corresponds to the case in which the predictor vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$ has real support $S=\{1,2,3,4,5\}$, meaning that only the first five variables affect the response.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ represents the error term, and each covariate $x_{ij}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(0,1),\ \mbox{for}\ j=1,\dots,p$.
Although only the first five variables are present in the response generating function $f(\mathbf{x}_{i})$, here considered the true support, the vector $\mathbf{x}_{i}$ ranges from $p=5$ to $p=100$ covariates, of which p5p-5 are irrelevant and represent the stochastic noise in the model.
@@ 2.2 Models Evaluated
Finally, the third group consists of hybrid Algorithms, in which each black-box model was fitted exclusively on the subset of variables previously selected by one of the regularized models, seeking to combine parsimony with predictive flexibility.
@@ 2.3 Simulations
This study evaluated the performance of regularized, pure, and hybrid black-box algorithms in twelve distinct scenarios, defined by combinations of five sample sizes (n = 50, 100, 200, 500, and 1,000) and four dimensionality levels (p = 5, 10, 50, and 100). For each scenario, ten simulated datasets were generated independently and replicated for each of the three proposed metrics (RMSE, Jaccard, and Recovery).
For each simulated dataset, 23 algorithms were applied: three regularized models (Ridge, Lasso, and Elastic Net), five pure black-box models (Random Forest, XGBoost, LightGBM, CatBoost, and H2O GBM), and 15 hybrid algorithms resulting from the combination of regularization-based and black-box algorithms.
Thus, each scenario involved $230$ runs $(10\text{ datasets}\times 23\text{ algorithms})$, replicated for 3 metrics and 20 associations of $n$ and $p$, totaling $20\times 230\times 3=13{,}800$ runs. Table 1 details the scenarios considered and the composition of the models evaluated by type. In each run, the data were randomly partitioned into training (80%) and testing (20%) sets to calculate the validation metrics.
@@
The regularized models then fitted using the penalty corresponding to the optimal value $\lambda_{\text{min}}$, obtained by cross-validation using the lowest RMSE criterion per feature subset. Based on the set of variables selected by these models, the hybrid algorithms then utilised the black-box methods based exclusively to the selected predictors. For comparison purposes, the pure black-box models were also fitted to the data, considering all available predictors, regardless of relevance. The final performance of the models was then evaluated based on the RMSE (Root Mean Squared Error), Jaccard index, and Recovery metrics.
@@
In Section  3, we focus on the values of $p=\{50\}$ and $n=\{200,500,1000\}$ and their main results for the purposes of brevity and clarity. However, the full range evaluated is presented in section 5, in table LABEL:tab:hybrid_models_full.
@@ 2.5 Computer Implementation
The results were saved in .csv and .RData files for reproducibility. Graphical visualizations were generated with the ggplot2 and viridis packages. The source code is publicly available [39].
@@ 3  Results
In this section, we present the results obtained in the simulation studies. The evaluation compared different modeling strategies, including regularized approaches, black-box models, and hybrid Algorithms that combine selection via regularization with nonlinear predictors. The evaluation metrics considered were the root mean square error (RMSE), the Jaccard index (for similarity between selected variables and the true set), and the recovery rate of relevant variables. The models in the subsections are analyzed segmented, allowing us to evaluate the relative performance of each class of approach under different complexity and generalization criteria.
@@ 3.1 Pure Black-Box Models
Table 2 presents the results of the main black-box models in their pure form, denoted as Full because they use the same algorithm for both variable selection and prediction, without combining them with regularized methods. Five algorithms are considered: CatBoost, H2O GBM, LightGBM, Random Forest, and XGBoost. The comparison includes three metrics at different sample sizes: root mean square error (RMSE), Jaccard index, and recovery rate.
Table 2: Pure black-box models with 90% noise ($p=50$). Values represent the mean (standard deviation) across 10 simulations. Bold highlights the best model per metric and sample size.
@@
RF RMSE 2.19 (0.73) 2.16 (0.05) 1.75 (0.52)
XGBoost RMSE 2.24 (0.15) 1.77 (0.53) 1.69 (0.05)
LightGBM RMSE 1.86 (0.56) 1.75 (0.05) 1.42 (0.42)
CatBoost RMSE 1.87 (0.09) 1.56 (0.04) 1.10 (0.48)
H2O GBM RMSE 2.02 (0.09) 1.80 (0.04) 1.45 (0.43)
RF Jaccard 0.83 (0.00) 0.83 (0.00) 0.83 (0.00)
XGBoost Jaccard 0.91 (0.12) 0.88 (0.15) 0.97 (0.07)
LightGBM Jaccard 0.76 (0.17) 0.76 (0.15) 0.89 (0.14)
CatBoost Jaccard 0.91 (0.12) 0.95 (0.08) 0.97 (0.07)
H2O GBM Jaccard 0.89 (0.16) 0.85 (0.15) 0.86 (0.18)
RF Recovery 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
XGBoost Recovery 0.98 (0.06) 1.00 (0.00) 1.00 (0.00)
LightGBM Recovery 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
CatBoost Recovery 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
H2O GBM Recovery 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
The results demonstrated that black-box models stand out for their high predictive power, with consistently low RMSE values, as already cited in the literature and therefore expected in this work. The LGBM model obtained the lowest RMSE ($1.86$) when taking a sample of size $200$ but its Jaccard index is relatively low compared to the others, obtaining the lowest metric for two of the three samples presented ($n=200$; $J=0.76$ and $n=500$; $J=0.76$). For larger samples ($n=500$ and $n=1000$) Catboost performed substantially better than the others, with root mean square error values of $1.56$ and $1.10$. The model also presents high Jaccard and Recovery values, even for smaller sample sizes ($n=200$ and $n=500$), indicating stability in identifying relevant variables. In terms of overlap with the true support, all models demonstrated excellent performance overall, with average recovery rates equal to or close to $1$ in all scenarios, with the exception of XGBoost, which recorded $0.98$ in smaller samples.